1. [4 pts) Formally prove (i.e. using € - 8 definition) that the function x3 is uniformly continuous on the interval (1,5).
Answer #1

solution- The function of is said to be uniformly continuous on s if and only if & Eri 7 f>o such that # By Es if I ae-y ! <fPeoof: Let Exo choose d=€ Then & x,y E (1,5] with le-yl cd we have Ifixe) – fcy)) = 1*3 – y3] = 1 (x-4) (x? + xy + y2) | :: =